# FlexVAR: Flexible Visual Autoregressive Modeling without Residual Prediction

### v1

@@ 6 Inference steps
7 Extent visual autoregressive modeling with Mamba
Unlike attention mechanisms that utilize explicit query-key-value (QKV) interactions to integrate context, Mamba faces challenges in handling bi-directional interaction. Therefore, prior Mamba-based visual autoregressive work [37] only used Mamba to model the unidirectional relationship between scales, relying on additional Transformer layers to process tokens within one scale.
In this work, we adopt a composition-recomposition strategy to obtain global information in Mamba network. Specifically, we utilize a Zigzag scanning strategy [19] over the spatial dimension. We alternate between eight distinct scanning paths across different Mamba layers (as shown in Fig. 12), which include:
@@ 7 Extent visual autoregressive modeling with Mamba
8 Qualitative results with different steps.
@@
9 Qualitative results with various resolutions.
[…]
10 Qualitative results with different VQVAE tokenizers.
@@
Generate images with GT prediction. We visualize the generated samples with VQVAE tokenizers from VAR, Llamagen, and ours, corresponding to the 2nd, 3rd and 5th results in Tab. 4. As shown in Fig. 16, the VAR tokenizer, trained with a residual paradigm, fails to generate images under GT prediction; the generation samples of Llamagen’s tokenizer are not up to the mark, due to its discrete tokens at intermediate steps being suboptimal.
11 Addtional Visual Results.
@@
[…]
Figure 14: Compared with different VAVAE tokenizers [44, 41] for multi-scale reconstructing images, we downsample the latent features in VQVAE to multiple scales and then use the VQVAE Decoder to reconstruct images. We upsample images $<$ 100 pixels using bilinear interpolation for a better view.
[…]
Figure 16: Some generated samples with different VQVAE tokenizers (Llamagen & VAR), corresponding to the 2nd and 3rd results in Tab. 4. We report the results with model scale -$d20$ trained 40 epochs ($\sim$ 70K iterations) on ImageNet-1K. Zoom in for a better view.
Figure 17: Some generated 256$\times$256 samples.

### v2

@@ Appendix A Inference steps
Appendix B Extent visual autoregressive modeling with Mamba
Unlike attention mechanisms that utilize explicit query-key-value (QKV) interactions to integrate context, Mamba faces challenges in handling bi-directional interaction. Therefore, prior Mamba-based visual autoregressive work ren2024m only used Mamba to model the unidirectional relationship between scales, relying on additional Transformer layers to process tokens within one scale.
In this work, we adopt a composition-recomposition strategy to obtain global information in Mamba network. Specifically, we utilize a Zigzag scanning strategy hu2024zigma over the spatial dimension. We alternate between eight distinct scanning paths across different Mamba layers (as shown in Fig. 11), which include:
@@ Appendix B Extent visual autoregressive modeling with Mamba
Appendix C Qualitative results with different steps.
@@
Appendix D Qualitative results with various resolutions.
Fig. 13 shows some generated samples with {256, 384, 512} resolutions. FlexVAR uses up to $\times256$ resolution images for training, it can generate images with higher resolutions such as 384 and 512. The generated images demonstrate strong semantic consistency across multiple scales, and the higher resolutions display more detailed clarity.
Appendix E Qualitative results with different VQVAE tokenizers.
@@
Generate images with GT prediction. We visualize the generated samples with VQVAE tokenizers from VAR, Llamagen, and ours, corresponding to the 2nd, 3rd and 5th results in Tab. 6 in the main paper. As shown in Fig. 15, the VAR tokenizer, trained with a residual paradigm, fails to generate images under GT prediction; the generation samples of Llamagen’s tokenizer are not up to the mark, due to its discrete tokens at intermediate steps being suboptimal.
Appendix F Additional Visual Results.
@@
Figure 12: Some generated samples with {6, 8, 10, 12} steps. Note the model is trained with steps $\leq$ 10. More steps typically result in better image details. Zoom in for a better view.
Figure 13: Some generated samples with {256, 384, 512} resolutions. Note the model is trained with a resolution of $\leq$ $\times256$. Zoom in for a better view.
Figure 14: Compared with different VQVAE tokenizers var ; llamagen for multi-scale reconstructing images, we downsample the latent features in VQVAE to multiple scales and then use the VQVAE Decoder to reconstruct images. We upsample images << 100 pixels using bilinear interpolation for a better view.
Figure 15: Some generated samples with different VQVAE tokenizers (Llamagen & VAR), corresponding to the 2nd and 3rd results in Tab. 6 in the main paper. We report the results with model scale -$d20$ trained 40 epochs ($\sim$ 70K iterations) on ImageNet-1K. Zoom in for a better view.
Figure 16: Some generated $\times256$ samples.